How do I find the vertex when it isn't in standard form and not a whole number?
f(x) = -4x^2-3x-5
I tried x of vertex = -b/2a
3/12
y of vertex I plugged 3/12 into the problem and then got lost.

Answer by stanbon(75887)

How do I find the vertex when it isn't in standard form and not a whole number?
f(x) = -4x^2-3x-5
I tried x of vertex = -b/2a = 3/(-8) = -3/8
----------------------
plug -3/8 into the problem to get (-4(-3/8)^2) -3(-3/8)-5 = -4.438..
Cheers,
Stan H.
===============